Use what you need of the following set of information to answer the question:

S0 = $1.1420/€            S0 = £0.7776/$                        S0 = $1.2860/£       

                        E (S1 (3-months)) = $1.1280/€                              F3-months = $1.1590/€

                 Where: € = euro; £ = British pound; E ( ) = Expected.

  1. If a currency dealer wants to use the above three-month forward rate to quote the three-month dollar as a percent premium or discount what rate would she quote? Show all calculations.

(Show all your calculations and explain your work. All currency amount answers should be presented with two decimal places, all currency rates with the same number of decimals they are given, and all rates of change and premiums/discounts with four decimal places)
Answer #1
Spot rate for $ in terms of Euro = 1/1.1420 = 0.8757
Forward rate for $ in terms of Euro = 1/1.1590 = 0.8628
The % premium/(discount) for $ in terms of Euro = 0.8628/0.8757 = -1.4731%
The above rate (discount) is for 3 months. Annualized rate =       -1.4731%*4 = -5.8924%
The rate of discount to be quoted is -5.8924% p.a.